You are sitting on a skateboard on the top of a hill. The hill has a 20% grade to it. The coefficient of rolling friction for you is μr “ 0.2. You have a mass of m “ 80kg. (a) What is the work done by gravity as when you roll a distance of 50m down the hill (that is along the hypotenuse of the hill, not elevation)? (b) What is the work done by rolling friction as when you roll a distance of 50m down the hill (that is along the hypotenuse of the hill, not elevation)? (c) What is the work done by normal force as when you roll a distance of 50m down the hill (that is along the hypotenuse of the hill, not elevation)?
Answer #1

Here, 20 % grade means the angle of incline is

= arctan ( 0.2)

= 11.3099 degree

-----------------------

(a) work done by gravity

W = force * distance * cos

W = mgcos ( 90 + 11.309) * 50

W = 80 * 9.8 * 50 * cos 101.3099

W = - 7687.75 J

------------------------------------

(b) work done by rolling friction

W = 80 * 9.8 * 50 * cos 180

W = - 39200 J

-------------------------------------

(c) work done by normal force

W = 80 * 9.8 * 50 * cos 90

W = 0 J